OpenStudy (anonymous):

show that lim x->0 [(2+x)sin(2+x) - 2sin2]/x = 2cos2+sin 2

OpenStudy (anonymous):

Let \[ f(x)= x \sin (x)\\ f'(x)=\sin (x)+x \cos (x) \] This is a hint. Compute f'(2) from the definition

OpenStudy (anonymous):

\[ \frac{f(x+2)-f(2)}{x}=\frac{(x +2) \sin (x+2)-2 \sin (2)}{x}\\ \lim_{x->0} \frac{f(x+2)-f(2)}{x}=f'(2)=\sin (2)+2 \cos (2) \]

OpenStudy (anonymous):

@eliassaab we cannot use L'hospital

OpenStudy (anonymous):

It is obvious that your limit is the derivative f'(2)

OpenStudy (anonymous):

Alternatively, assuming we don't know how to find derivatives yet, you can use the angle sum formula for sine: \[\sin(x+2)=\sin x\cos2+\cos x\sin 2\] So, \[\lim_{x\to0}\frac{(x+2)\sin(x+2)-2\sin 2}{x}=\lim_{x\to0}\frac{(x+2)(\sin x\cos2+\cos x\sin2)-2\sin 2}{x}\\ ~~~~~~~~=\lim_{x\to0}\frac{x\sin x\cos2+2\sin x\cos2+x\cos x\sin2+2\cos x\sin2-2\sin2}{x}\\ ~~~~~~~~=\cos2\left(\lim_{x\to0}\frac{\sin x(x+2)}{x}\right)+\sin2\left(\lim_{x\to0}\frac{x\cos x}{x}+2\lim_{x\to0}\frac{\cos x-1}{x}\right) \] which can be reduced if you know that \[\lim_{x\to0}\frac{\sin ax}{ax}=1~~~\text{and}~~~\lim_{x\to0}\frac{1-\cos ax}{ax}=0~~~\text{for }a\not=0\]

OpenStudy (anonymous):

@SithsAndGiggles , your method is correct. It is longer and you have to find the derivative anyway.

OpenStudy (anonymous):

@SithsAndGiggles how do we show the last step ie: 2cos2+sin 2

OpenStudy (anonymous):

ok i got it @SithsAndGiggles

OpenStudy (anonymous):

Thanks for your help !!

OpenStudy (anonymous):

\[\cos2\left(\lim_{x\to0}\frac{\sin x(x+2)}{x}\right)+\sin2\left(\lim_{x\to0}\frac{x\cos x}{x}+2\lim_{x\to0}\frac{\cos x-1}{x}\right)\] One of the properties of limits you should be aware of is that the limit of a product is the same as the product of the limits; mathematically, you have \[\lim_{x\to c}f(x)g(x)=\left(\lim_{x\to c}f(x)\right)\left(\lim_{x\to c}g(x)\right)\] You can apply this property directly to the first limit: \[\cos2\left(\lim_{x\to0}\frac{\sin x(x+2)}{x}\right)=\cos2\left(\color{red}{\lim_{x\to0}\frac{\sin x}{x}}\right)\left(\lim_{x\to0}(x+2)\right)\] In the second limit, the \(\dfrac{x}{x}\) just becomes 1, so you're left with \(\cos x\). The third limit is known.

OpenStudy (anonymous):

You're welcome!

OpenStudy (anonymous):

Here is a point that many people do not know. To prove that the derivative of sin(x) is cos(x), we use the geometric fact that sin(x)/x approaches 1 when x approaches zero. You really cannot use L'hospital rule to show that that sin(x)/x approaches 1 when x approaches zero, but people do. C'est la vie.